# Existence for Stable Rotating Star-Planet Systems

### v1

@@ 5 Bound of the Supports of Density Functions
In fact, the convergence up to a subsequence can be strengthened to convergence of the whole sequence. Since the non-rotating minimizer is unique up to translation due to Theorem 3.6, we get $\left\{\widetilde{\rho_{m_{k}}}\right\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ up to translations. By the same arguments, we know for any subsequence $\left\{\widetilde{\rho_{m_{k}}}\right\}$, we can always pick a subsequence $\left\{\widetilde{\rho_{m_{k_{l}}}}\right\}$ of $\left\{\widetilde{\rho_{m_{k}}}\right\}$ which converges strongly to $\sigma$ up to translations. Thus, by an argument of contradiction, we get, up to translations, $\left\{\widetilde{\rho_{m}}\right\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$. ∎
@@
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{1-m}}=A(1-m)\rho_{1-m}\left({B(1-m)x}\right)$ as in Definition 5.1, we have, up to translations, $\{\widetilde{\rho_{1-m}}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
@@
Thus ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\widetilde{\rho_{1-m}}\right)}}=e_{0}<0$, and we can follow the similar proof as in Theorem 5.9 above to get the result. ∎
@@
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, we have, up to translations, $\{\rho_{1-m}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
@@
Thus ${\lim\limits_{m\rightarrow 0}{E_{0}\left({\rho_{1-m}}\right)}}=e_{0}<0$, and we can follow the similar proof as in Theorem 5.9 above to get the result. ∎
@@
Then we have ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\rho_{m}\right)}}=0$, which violates the condition of concentration-compactness lemma [42, Theorem II.2 and Corollary II.1]. In this case, we can actually show that $\rho_{m}$ converges to 0: from the scaling relation (36) and the fact $\lim\limits_{m\rightarrow 0}{A(m)}$$\infty$ and ${\lim\limits_{m\rightarrow 0}{B(m)}}=0$, we can show $\left.\left\|\rho_{m}\right\|_{L^{\frac{4}{3}}}=\left({A(m)}\right)^{-\frac{4}{3}}\left({B(m)}\right)^{3}\left\|\widetilde{\rho_{m}}\right\|_{L^{\frac{4}{3}}}\rightarrow 0\right$. as $m\rightarrow 0$. However, this vanishing limit result is not enough. We seek a stronger estimate that will allow us to estimate the Lagrange multiplier and then the support size of $\rho_{m}$ using the arguments in the following. This is the reason we introduce the scaling density.
@@
Note the convergence holds true for the whole sequence instead of up to subsequence.
@@ 5.2 Uniform bound of scaling densities in $L^{\infty}\left(\mathbb{R}^{3}\right)$
Thanks to the convergence results in subsection 5.1, now we can estimate the uniform $L^{\infty}$ bound of scaling densities $\left\{\widetilde{\rho_{m}}\right\}$ when $m$ is sufficiently small. Since $\widetilde{\rho_{m}}\rightarrow\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ (Theorem 5.9), thus when $m$ is sufficiently small, $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$. The next steps are the following: (1) $\left\{V_{\widetilde{\rho_{m}}}\right\}$ are bounded uniformly in $L^{12}\left(\mathbb{R}^{3}\right)$; (2) $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{6}\left(\mathbb{R}^{3}\right)$; (3) $\left\{V_{\widetilde{\rho_{m}}}\right\}$ are bounded uniformly in $L^{\infty}\left(\mathbb{R}^{3}\right)$; (4) $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{\infty}\left(\mathbb{R}^{3}\right)$. Since we consider $\left\{\widetilde{\rho_{m}}\right\}$ and $\left\{V_{\widetilde{\rho_{m}}}\right\}$ alternatively, we call such method the bootstrap method.
@@
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{m}}=A(m)\rho_{m}\left({B(m)x}\right)$ as in Definition 5.1, then $\exists C>0$, $\delta>0$, $\forall 0<m<\delta$, we have $\widetilde{\rho_{m}}\in L^{\infty}\left(\mathbb{R}^{3}\right)$, and $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq C$, $V_{\widetilde{\rho_{m}}}\in L^{\infty}\left(\mathbb{R}^{3}\right)$, and $\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq C$, where $V_{\widetilde{\rho_{m}}}(x)={\int{\frac{\widetilde{\rho_{m}}(y)}{\left|{x-y}\right|}d^{3}y}}$ is the potential from $\widetilde{\rho_{m}}$. Moreover, $V_{\widetilde{\rho_{m}}}(x)$ is continuous.
@@
$\left\|\widetilde{\rho_{m}}\right\|_{L^{12{({\gamma-1})}}}=\left({\int{{\widetilde{\rho_{m}}}^{12{({\gamma-1})}}(x)d^{3}x}}\right)^{\frac{1}{12{({\gamma-1})}}}\leq{\widetilde{C}\left({\int{V_{\widetilde{\rho_{m}}}^{12}(x)d^{3}x}}\right)}^{\frac{1}{12{({\gamma-1})}}}=\widetilde{C}\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{12}}^{\frac{1}{\gamma-1}}\leq C_{2}$
By assumption $\gamma>\frac{3}{2}>\frac{4}{3}$, together with Interpolation Inequality Proposition A.2 we have $\left\|\widetilde{\rho_{m}}\right\|_{L^{12}}\leq\left\|\widetilde{\rho_{m}}\right\|_{L^{12{({\gamma-1})}}}^{\alpha}\cdot\left\|\widetilde{\rho_{m}}\right\|_{L^{1}}^{1-\alpha}\leq C$ for some $C>0$. Since we already know $\left\|\widetilde{\rho_{m}}\right\|_{L^{1}{(\mathbb{R}^{3})}}=1$ and $\left\|\widetilde{\rho_{m}}\right\|_{L^{12}{(\mathbb{R}^{3})}}$ are bounded uniformly, thanks to Proposition A.4, we have $V_{\widetilde{\rho_{m}}}$ is continuous and $\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{\infty}{(\mathbb{R}^{3})}}$ is bounded uniformly. Again by ${\widetilde{\rho_{m}}}^{\gamma-1}\leq CV_{\widetilde{\rho_{m}}}$ we know $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}$ is also bounded uniformly. ∎
@@
The uniform bound in $L^{\infty}(\mathbb{R}^{3})$ can be comparable with Li’s result [38, Proposition 1.4], where the rotating stars (corresponding to single-star case) are considered and the bound $C(m)$ however depends on $m$. Although one might conjecture $C(m)$ is continuous when $m>0$, they may still need to consider if $\lim\limits_{m\rightarrow 0}{\sup{C(m)}}$$\infty$ in order to obtain a result analogous to Theorem 5.14.
@@
By the similar arguments, we can obtain that when $m$ is sufficient small, $\left\{\widetilde{\rho_{1-m}}\right\}$ and $\left\{\rho_{1-m}\right\}$ are also bounded uniformly in $L^{\infty}(\mathbb{R}^{3})$.
@@
Similarly as Remark 3.12, once we know $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}$ is bounded uniformly, by definition we have $\rho_{m}=\frac{1}{A}\widetilde{\rho_{m}}\left(\frac{x}{B}\right)$, $A=m^{-\frac{2}{3\gamma-4}}$, thus ${\lim\limits_{m\rightarrow 0}\left\|\rho_{m}\right\|_{L^{\infty}(\mathbb{R}^{3})}}=0$.
@@ 5.3 Estimate for the supports of densities
In this subsection, we finish the estimate for the supports of scaling densities as well as the original densities. We first seek an upper bound for the Lagrange multiplier. To avoid confusion with the Lagrange multiplier $\lambda_{m}$ corresponding to star planet system, we will use $\kappa_{m}$ to represent the corresponding non-rotating Lagrange multiplier with mass $m$ in this subsection.
@@
Similar to the arguments in the proof of Theorem 4.5 (see also [4, Lemma 2] [15, Proposition 5.8] [16, Lemma 2.21]), it can be shown that given a constrained minimizer $\rho(m)$, then $\rho(m)$ satisfies Euler-Lagrange equation in the following sense:
@@
where $\mathcal{A}(\rho)(x)=\frac{K\left(\rho^{\gamma}(x)\right)}{\gamma-1},\mathcal{A}^{\prime}(\rho)(x)=\frac{K\gamma\left(\rho^{\gamma-1}(x)\right)}{\gamma-1}$. Due to (30) (31), we have $\mathcal{A}^{\prime}\left(\rho_{m}\right)(x)=\mathcal{A}^{\prime}\left(\frac{1}{A}\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)$, $V_{\rho_{m}}(x)=\frac{B^{2}}{A}V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)$. Therefore, we have
@@
By Theorem 3.6 (v) and the fact $\sigma$ has mass 1, we know $0<\mu\left(\left\{x\in\mathbb{R}^{3}\mid\sigma(x)>0\right\}\right)<\infty$. Since $0\leq\sigma<C$ a.e. on $\mathbb{R}^{3}$, let $E_{n}=\left\{x\in\mathbb{R}^{3}\mid\frac{C}{n}\leq\sigma(x)<C\right\}$, then $\mathbb{R}^{3}=\cup_{n=1}^{\infty}E_{n}$. And then $1=\int_{\mathbb{R}^{3}}\sigma=\sum_{n=1}^{\infty}\int_{E_{n}}\sigma$. Then $\exists N>0$, such that $\frac{1}{2}<\sum_{n=1}^{N}\int_{E_{n}}\sigma<\sum_{n=1}^{N}\int_{E_{n}}C=C\mu(\{\{x\in\mathbb{R}^{3}\mid\frac{C}{N}\leq\sigma(x)<C\})$. Therefore, let $M:=\mu\left(\left\{x\in\mathbb{R}^{3}\mid\frac{C}{N}<\sigma(x)<C\right\}\right)$, then $0<\frac{1}{2C}<M<\infty$. Thus $\exists\delta>0,\forall 0<m<\delta$, due to the claim we have
@@
Note that the above proof relies on convergence in measure. In McCann’s paper [46, Lemma 6.5], only almost everywhere convergence up to subsequence is used. Nonetheless, since in this paper we assume $m$ is very small, relying solely on almost everywhere convergence up to subsequence will lead to the following two issues. One is that it is up to a subsequence, i.e. $\widetilde{\rho_{m_{k}}}(x)\rightarrow\sigma(x)$ a.e. on $\mathbb{R}^{3}$ instead of $\left\{\widetilde{\rho_{m}}\right\}$ itself, while we hope the bound for the Lagrange multiplier should be valid for all sufficient small mass. Another one is when $m\rightarrow 0$, not only $\widetilde{\rho_{m}}(x)\rightarrow\sigma(x)$ a.e. on $\mathbb{R}^{3}$, but the measures of relevant sets above go to 0. For example, when $m\rightarrow 0$, $\mu\left(\left\{x\in\mathbb{R}^{3}\mid\sigma\left(\frac{x}{B(m)}\right)>0\right\}\right)=B^{3}(m)\mu\left(\left\{x\in\mathbb{R}^{3}\mid\sigma(x)>0\right\}\right)\rightarrow 0$. Therefore, we need to rule out the case that $\widetilde{E}$ is empty; otherwise, it is not clear whether there exists an $x$ that satisfies the last inequality in (41). Therefore, we make more use of $L^{p}$ convergence—namely, that it implies convergence in measure, to solve the issues.
Before we finally estimate the sizes of supports of scaling densities, we supplement the proof of a proposition about the $L^{\infty}$ bound of potential mentioned in [46, Proposition 6.6], which can be viewed as a complement to [4, Proposition 5], in which the indices $\frac{2}{3}$ and $\frac{1}{3}$ can be approximated but cannot be reached.
Proposition 5.20 (Bound of Potential in $L^{\infty}\left(\mathbb{R}^{3}\right)$ [46, Section 6]).
$\exists k>0,\forall\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, one has $\left\|V_{\rho}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq k\|\rho\|_{L^{1}\left(\mathbb{R}^{3}\right)}^{\frac{2}{3}}\|\rho\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}^{\frac{1}{3}}$, where $V_{\rho}(x)=\int\frac{\rho(y)}{|x-y|}d^{3}y$ is the potential from $\rho$.
@@
$V_{\rho}(x)$ $=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}d^{3}y$
@@
Take $m$ and $\epsilon$ small enough so that constrained minimizers $\rho(m)$ exist by Theorem 4.14 and $A(m)^{\gamma-1}\cdot\lambda_{m}\leq\kappa_{1}+\epsilon<$ 0 by Lemma 5.18, and the velocity $v(x,m)$ satisfies $v(x,m)^{2}\leq Cm^{2}$ by Corollary 5.5. By the construction of $\Omega_{m}$ and $\Omega_{1-m}$, we have $\operatorname{dist}\left(\Omega_{m},\Omega_{1-m}\right)=\frac{\eta}{2}$, $\eta=\frac{J^{2}}{\mu^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$, thus $\exists C>0,\forall x\in\Omega_{m}$, $V_{\rho_{1-m}}(x)<Cm^{2}$. In the Euler-Lagrange equation (ELm), we know for almost every $x\in\Omega_{m}$,

### v2

@@ 5 Bound of the Supports of Density Functions
In fact, the convergence up to a subsequence can be strengthened to the convergence of the whole sequence. Since the non-rotating minimizer is unique up to translation due to Theorem 3.6, we get $\left\{\widetilde{\rho_{m_{k}}}\right\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ up to translations. By the same arguments, we know for any subsequence $\left\{\widetilde{\rho_{m_{k}}}\right\}$, we can always pick a subsequence $\left\{\widetilde{\rho_{m_{k_{l}}}}\right\}$ of $\left\{\widetilde{\rho_{m_{k}}}\right\}$ which converges strongly to $\sigma$ up to translations. Thus, by an argument of contradiction, we get, up to translations, $\left\{\widetilde{\rho_{m}}\right\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$. ∎
@@
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$, and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{1-m}}=A(1-m)\rho_{1-m}\left({B(1-m)x}\right)$ as in Definition 5.1, we have, up to translations, $\{\widetilde{\rho_{1-m}}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
@@
Thus ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\widetilde{\rho_{1-m}}\right)}}=e_{0}<0$, and we can follow a proof similar to that in Theorem 5.9 above to get the result. ∎
@@
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$, and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, we have, up to translations, $\{\rho_{1-m}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
@@
Thus ${\lim\limits_{m\rightarrow 0}{E_{0}\left({\rho_{1-m}}\right)}}=e_{0}<0$, and we can follow a proof similar to that in Theorem 5.9 above to get the result. ∎
@@
Then we have ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\rho_{m}\right)}}=0$, which violates the condition of concentration-compactness lemma [22, Theorem II.2 and Corollary II.1]. In this case, we can actually show that $\rho_{m}$ converges to 0: from the scaling relation (37) and the fact $\lim\limits_{m\rightarrow 0}{A(m)}$$\infty$ and ${\lim\limits_{m\rightarrow 0}{B(m)}}=0$, we can show $\left.\left\|\rho_{m}\right\|_{L^{\frac{4}{3}}}=\left({A(m)}\right)^{-\frac{4}{3}}\left({B(m)}\right)^{3}\left\|\widetilde{\rho_{m}}\right\|_{L^{\frac{4}{3}}}\rightarrow 0\right$. as $m\rightarrow 0$. However, this vanishing limit result is not enough. We seek a stronger estimate that will allow us to estimate the Lagrange multiplier and then the support size of $\rho_{m}$, as we will see later. This is the reason we introduce the scaling density.
@@
Note the convergence holds true for the whole sequence instead of up to a subsequence.
@@ 5.2 Uniform bound of scaling densities in $L^{\infty}\left(\mathbb{R}^{3}\right)$
Thanks to the convergence results in subsection 5.1, now we can estimate the uniform $L^{\infty}$ bound of scaling densities $\left\{\widetilde{\rho_{m}}\right\}$ when $m$ is sufficiently small. Since $\widetilde{\rho_{m}}\rightarrow\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ (Theorem 5.9), we know $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ when $m$ is sufficiently small. The next steps are to apply bootstrap method as follows: (1) $\left\{V_{\widetilde{\rho_{m}}}\right\}$ are bounded uniformly in $L^{12}\left(\mathbb{R}^{3}\right)$; (2) $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{6}\left(\mathbb{R}^{3}\right)$; (3) $\left\{V_{\widetilde{\rho_{m}}}\right\}$ are bounded uniformly in $L^{\infty}\left(\mathbb{R}^{3}\right)$; (4) $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{\infty}\left(\mathbb{R}^{3}\right)$.
@@
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{m}}=A(m)\rho_{m}\left({B(m)x}\right)$ as in Definition 5.1, then $\exists C>0$, $\delta>0$, $\forall 0<m<\delta$, we have $\widetilde{\rho_{m}}\in L^{\infty}\left(\mathbb{R}^{3}\right)$, and $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq C$, $V_{\widetilde{\rho_{m}}}\in L^{\infty}\left(\mathbb{R}^{3}\right)$, and $\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq C$, where $V_{\widetilde{\rho_{m}}}(x)={\int_{\mathbb{R}^{3}}{\frac{\widetilde{\rho_{m}}(y)}{\left|{x-y}\right|}\,dy}}$ is the potential from $\widetilde{\rho_{m}}$. Moreover, $V_{\widetilde{\rho_{m}}}(x)$ is continuous.
@@
$\left\|\widetilde{\rho_{m}}\right\|_{L^{12{({\gamma-1})}}}=\left({\int_{\mathbb{R}^{3}}{{\widetilde{\rho_{m}}}^{12{({\gamma-1})}}(x)\,dx}}\right)^{\frac{1}{12{({\gamma-1})}}}\leq{\widetilde{C}\left({\int_{\mathbb{R}^{3}}{V_{\widetilde{\rho_{m}}}^{12}(x)\,dx}}\right)}^{\frac{1}{12{({\gamma-1})}}}=\widetilde{C}\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{12}}^{\frac{1}{\gamma-1}}\leq C_{2}$
By assumption $\gamma>\frac{3}{2}>\frac{4}{3}$, together with Interpolation Inequality [6, Section 4.2] we have $\left\|\widetilde{\rho_{m}}\right\|_{L^{12}}\leq\left\|\widetilde{\rho_{m}}\right\|_{L^{12{({\gamma-1})}}}^{\alpha}\cdot\left\|\widetilde{\rho_{m}}\right\|_{L^{1}}^{1-\alpha}\leq C$ for some $C>0$. Since we already know $\left\|\widetilde{\rho_{m}}\right\|_{L^{1}{(\mathbb{R}^{3})}}=1$ and $\left\|\widetilde{\rho_{m}}\right\|_{L^{12}{(\mathbb{R}^{3})}}$ are bounded uniformly, due to Proposition A.2, we have $V_{\widetilde{\rho_{m}}}$ is continuous and $\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{\infty}{(\mathbb{R}^{3})}}$ is bounded uniformly. Again by ${\widetilde{\rho_{m}}}^{\gamma-1}\leq CV_{\widetilde{\rho_{m}}}$ we know $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}$ is also bounded uniformly. ∎
@@
The uniform bound in $L^{\infty}(\mathbb{R}^{3})$ can be comparable with Li’s result [18, Proposition 1.4], where the rotating single stars are considered and the bound $C(m)$ however depends on $m$. Hence, if it can be shown that $\lim\limits_{m\rightarrow 0}{\sup{C(m)}}$$\infty$, then an analogous result to Theorem 5.14 should also hold.
@@
By the similar arguments, we can obtain that when $m$ is sufficiently small, $\left\{\widetilde{\rho_{1-m}}\right\}$ and $\left\{\rho_{1-m}\right\}$ are also bounded uniformly in $L^{\infty}(\mathbb{R}^{3})$.
@@
Similar to the discussion in Remark 3.10, once we know $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}$ is bounded uniformly, by definition we have $\rho_{m}=\frac{1}{A}\widetilde{\rho_{m}}\left(\frac{x}{B}\right)$, $A=m^{-\frac{2}{3\gamma-4}}$, thus ${\lim\limits_{m\rightarrow 0}\left\|\rho_{m}\right\|_{L^{\infty}(\mathbb{R}^{3})}}=0$.
@@ 5.3 Estimate for the supports of densities
In this subsection, we finish the estimate for the supports of scaling densities as well as the original densities. We first seek an upper bound for the Lagrange multiplier. To avoid confusion with the Lagrange multiplier $\lambda_{m}$ corresponding to star-planet system, we will use $\kappa_{m}$ to represent the corresponding non-rotating Lagrange multiplier with mass $m$ in this subsection.
@@
Similar to the arguments in the proof of Theorem 4.5 (see also [4, Lemma 2] [10, Section 5] [11, Section 2]), it can be shown that given a constrained minimizer $\rho(m)$, $\rho(m)$ satisfies Euler-Lagrange equation in the following sense:
@@
where $\mathcal{A}(\rho)(x)=\frac{K\left(\rho^{\gamma}(x)\right)}{\gamma-1}$, $\mathcal{A}^{\prime}(\rho)(x)=\frac{K\gamma\left(\rho^{\gamma-1}(x)\right)}{\gamma-1}$. Due to (31) (32), we have
$\mathcal{A}^{\prime}\left(\rho_{m}\right)(x)=\mathcal{A}^{\prime}\left(\frac{1}{A}\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)$
$V_{\rho_{m}}(x)=\frac{B^{2}}{A}V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)$
Therefore, we have
@@
By Theorem 3.6 (v) and the fact $\sigma$ has mass 1, we know $0<\mu\left(\left\{x\in\mathbb{R}^{3}\mid\sigma(x)>0\right\}\right)<\infty$. Since $0\leq\sigma<C$ a.e. on $\mathbb{R}^{3}$, let $E_{n}=\left\{x\in\mathbb{R}^{3}\mid\frac{C}{n}\leq\sigma(x)<C\right\}$, then $\mathbb{R}^{3}=\cup_{n=1}^{\infty}E_{n}$. And then $1=\int_{\mathbb{R}^{3}}\sigma\,dx=\sum_{n=1}^{\infty}\int_{E_{n}}\sigma\,dx$. Then $\exists N>0$, such that
$\frac{1}{2}<\sum_{n=1}^{N}\int_{E_{n}}\sigma\,dx<\sum_{n=1}^{N}\int_{E_{n}}C\,dx=C\mu(\{\{x\in\mathbb{R}^{3}\mid\frac{C}{N}\leq\sigma(x)<C\})$
Therefore, let $M:=\mu\left(\left\{x\in\mathbb{R}^{3}\mid\frac{C}{N}<\sigma(x)<C\right\}\right)$, then $0<\frac{1}{2C}<M<\infty$. Thus $\exists\delta>0$, $\forall 0<m<\delta$, due to the claim we have
@@
$\widetilde{\rho_{m}}\left(\frac{x}{B}\right)\geq\sigma\left(\frac{x}{B}\right)-\left|\widetilde{\rho_{m}}\left(\frac{x}{B}\right)-\sigma\left(\frac{x}{B}\right)\right|>\frac{C}{N}-\frac{C}{3N}>0$
Hence $\rho_{m}(x)=\frac{1}{A}\widetilde{\rho_{m}}\left(\frac{x}{B}\right)>0$. Thus $x\in\Omega_{m}$ (up to translations),
@@
Note that the above proof relies on convergence in measure. In McCann’s paper [24, Lemma 6.5], only almost everywhere convergence up to a subsequence is used. Nonetheless, since in this paper we assume $m$ is very small, relying solely on almost everywhere convergence up to a subsequence will lead to the following two issues. One is that it is up to a subsequence, i.e., $\widetilde{\rho_{m_{k}}}(x)\rightarrow\sigma(x)$ a.e. on $\mathbb{R}^{3}$ instead of $\left\{\widetilde{\rho_{m}}\right\}$ itself, while we hope the bound for the Lagrange multiplier should be valid for all sufficiently small mass. Another one is when $m\rightarrow 0$, not only $\widetilde{\rho_{m}}(x)\rightarrow\sigma(x)$ a.e. on $\mathbb{R}^{3}$, but the measures of relevant sets above go to 0. For example, when $m\rightarrow 0$, $\mu\left(\left\{x\in\mathbb{R}^{3}\mid\sigma\left(\frac{x}{B(m)}\right)>0\right\}\right)=B^{3}(m)\mu\left(\left\{x\in\mathbb{R}^{3}\mid\sigma(x)>0\right\}\right)\rightarrow 0$. Therefore, we need to rule out the case that $\widetilde{E}$ is empty; otherwise, it is not clear whether there exists an $x$ that satisfies the last inequality in (42). Therefore, we make more use of $L^{p}$ convergence—namely, that it implies convergence in measure, to solve the issues.
Before we finally estimate the sizes of supports of scaling densities, we supplement the proof of a proposition about the $L^{\infty}$ bound of potential mentioned in [24, Proposition 6.6], which can be viewed as a complement to [4, Proposition 5], in which the indices $\frac{2}{3}$ and $\frac{1}{3}$ can be approximated but cannot be reached.
Proposition 5.20 (Bound of Potential in $L^{\infty}\left(\mathbb{R}^{3}\right)$ [24, Section 6]).
$\exists k>0,\forall\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, one has $\left\|V_{\rho}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq k\|\rho\|_{L^{1}\left(\mathbb{R}^{3}\right)}^{\frac{2}{3}}\|\rho\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}^{\frac{1}{3}}$, where $V_{\rho}(x)=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}\,dy$ is the potential from $\rho$.
@@
Let $B_{R}(\cdot)$ be the open ball defined by $B_{R}(\cdot)\coloneq\left\{x\in\mathbb{R}^{3}\mid|x-\cdot|<R\right\}$.
$V_{\rho}(x)$ $=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}\,dy$
@@
Take $m$ and $\epsilon$ small enough so that constrained minimizers $\rho(m)$ exist by Theorem 4.14 and $A(m)^{\gamma-1}\cdot\lambda_{m}\leq\kappa_{1}+\epsilon<$ 0 by Lemma 5.18, and the velocity $v(x,m)$ satisfies $v(x,m)^{2}\leq Cm^{2}$ by Corollary 5.5. By the construction of $\Omega_{m}$ and $\Omega_{1-m}$, we have $\operatorname{dist}\left(\Omega_{m},\Omega_{1-m}\right)=\frac{\eta}{2}$, where $\eta=\frac{J^{2}}{\mu_{r}^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$, thus $\exists C>0$, $\forall x\in\Omega_{m}$, $V_{\rho_{1-m}}(x)<Cm^{2}$. In the Euler-Lagrange equation (ELm), we know for almost every $x\in\Omega_{m}$,